OpenStudy (kkrpbra):

Adele is watching a cartoon on a rectangular TV screen. The screen is 28 inches wide and 16 inches high. What is the perimeter of the TV screen in units of inches? Choose the correct expression.

OpenStudy (zepp):

 \(p = H+W+H+W~~or~~ p= 2(H+W)\)